# Astrocyte Regulated Neuromorphic Central Pattern Generator Control of Legged Robotic Locomotion

### v2

@@ III-B2 Inter-limb Connection
Recent works have elaborated that there are excitatory and inhibitory synapses from the motor neurons for one limb to the motor neurons for another limb [10]. In this paper, the phase difference among the limbs is modulated by the inter-limb excitatory/inhibitory connections. As shown in Figure 1d, the thigh muscles from one limb send axons to all the muscles in all the other limbs. For example, the motor neurons of the thigh extensor in the front-left limb send connections to the thigh flexors and the extensors in front-right, rear-left and rear-right limbs. The calf motor neurons are not involved in the inter-limb connection. The source motor neuron pool and the target motor neuron pool are fully connected, but the entire connection shares a single synapse strength value. Such a connection could be expressed as a table of size 8 by 8, shown in Figure 2a, where the value in row $i$ and column $j$ represents the synapse weight of the connection projected from muscle $i$ to muscle $j$. The inter-limb connection is trained by reward-modulated Spike Timing Dependent Plasticity (STDP), including the necessary modulation of astrocytes, which will be introduced in a later section.
@@ IV Gait Learning Algorithm
IV-A Reward-Modulated STDP for Gait Search
Previous works have discussed the reward-modulated STDP for encouraging actions through delayed reward signal [18]. The basic idea of gait search is to reward the gait, i.e., the phase difference of the bursts among all the motor neuron pools, providing a larger forward speed and more stable torso attitude. The effect of different collaboration of the muscles is usually delayed, i.e., a good gait leads to a large forward speed, but the speeding up procedure still costs a duration of time. Hence, reward-modulated STDP is necessary for gait search. The reward function is given by:
@@ IV-A Reward-Modulated STDP for Gait Search
where, $\mathrm{STDP}_{x,y}(t)$ is the STDP signal from motor neuron pool $x$ to motor neuron pool $y$. $\tau_{\mathrm{STDP}}$ is the STDP signal decay time constant. $t_{y\mathrm{\_spike}}$ and $t_{x\mathrm{\_spike}}$ are the times of firing of motor neuron pool $y$ and $x$. $u_{x}(t)$ and $u_{y}(t)$ are the spike traces of motor neuron pool $x$ and $y$ at time $t$. $\tau_{\mathrm{trace}}$ is the spike trace decay time constant. $\eta_{\mathrm{negative\_relative}}$ is the relative negative learning rate, which controls the contribution of negative STDP signal in learning. $\eta_{\mathrm{eff}}(t)$ is the effective learning rate, which is the product of training progress $\mathrm{Progress}(t)$ and learning rate $\eta$. The definition of training progress can be found in the Methods section. $r_{\mathrm{eff}}(t)$ is the effective reward for STDP learning. $\overline{r_{(t-0.1s,t)}}$ is the sliding window average of the reward function from 0.1 seconds before to the current time. $c_{\mathrm{average}}$ is the averaged reward coefficient. By subtracting a fraction of $\overline{r_{(t-0.1s,t)}}$, the system encourages the concurrent firing of the muscles which provides higher reward than the average. It also discourages the concurrent firing of the muscles leading to the falling of the robot, at which time the angular velocities of the torso are large. $\zeta_{w_{x,y}}(t)$ is a weight constraint factor preventing the weight to be out of the range of $[w_{\mathrm{min}},w_{\mathrm{max}}]$.
@@
It is worth noting here that the STDP learning algorithm is usually applied in situations where excitatory synapses play a significant role [7]. The learning rate for post-synaptic firing is usually much larger than its counterpart for pre-synaptic firing. As a result, the learning of inhibitory synapses, which are critical in many neural systems like ours, is largely unexplored. As we show next, the astrocyte regulation of synaptic plasticity is found to be critical for building inhibitory connections in the CPG.
IV-B Astrocyte Regulation
It is reported that Endocannabinoid receptor $\mathrm{CB}_{1}$ is expressed by the astrocytes in multiple brain regions and in the dorsal horn of the spinal cord [35]. In this paper, we regard 2-arachidonyl glycerol (2-AG), a type of endocannabinoid, as the effective neurotransmitter signaling from motor neurons to the astrocytes. The motor neurons release a constant amount of 2-AG every time when there is an action potential. The concentration of 2-AG decays exponentially when no motor neuron fires:
Li-Rinzel model [23] can be used to describe the cytoplasmic dynamics of $\mathrm{IP}_{3}$ and $\mathrm{Ca}^{2+}$ in the astrocytes. The activation of $\mathrm{CB}_{1}$ receptor results in $\mathrm{IP}_{3}$ release in the astrocyte cytoplasm. $\mathrm{IP}_{3}$ further regulates the release of $\mathrm{Ca}^{2+}$ from the Endoplasmic Reticulum (ER) [42]. The transportation of $\mathrm{Ca}^{2+}$ between the cytoplasm and ER can be divided into three components: the channel current from ER to the cytoplasm, $J_{\mathrm{chan}}$; the passive leakage current from the ER to the cytoplasm, $J_{\mathrm{leak}}$; the active pumping current from the cytoplasm to ER, $J_{\mathrm{pump}}$. In summary:
The $[\mathrm{Ca}^{2+}]$ here is the concentration of cytoplasmic $\mathrm{Ca}^{2+}$. Readers seeking a detailed computational model are directed to [42].
As no past work provides a computational model of the ATP release from astrocytes in the spinal cord (to the best of our knowledge), we assume that a fixed amount of ATP is released (with a given refractory period) when the concentration of $\mathrm{Ca}^{2+}$ is higher than a given threshold, which is similar to the release of glutamate in the hippocampus. Additionally, as the ATP is always hydrolyzed into adenosine in a very short time period, adenosine is directly adopted as the effective output from the astrocyte:
Here, $[\mathrm{ADO}]$ is the concentration of adenosine in the nearby region of the astrocyte, $\tau_{\mathrm{ADO}}$ is the adenosine decay time constant, $r_{\mathrm{ADO}}$ is the production rate of adenosine, $t_{\mathrm{Ca}^{2+},\mathrm{ADO\_prev}}$ is the time instant when $[\mathrm{Ca}^{2+}]$ is above the threshold satisfying the condition that the time from the last adenosine release is not shorter than the adenosine refractory period. The adenosine received by the motor neurons negatively contributes to the synapse strength sent to this motor neuron pool. Figure 2b illustrates the negative feedback relationship between the motor neuron pool and the adenosine. Therefore, at every time step, astrocyte-induced weight update from motor neuron pool $x$ to motor neuron pool $y$ is:
@@ IV-B Astrocyte Regulation
where, $\eta_{\mathrm{ADO}}$ is the efficacy of adenosine in building inhibitory synapses, $\mathrm{Progress}(t)$ is the training progress, $[\mathrm{ADO}_{y}]$ is the adenosine concentration at motor neuron pool $y$, $\zeta_{w_{x,y}}(t)$ is the weight constraint factor. The complete weight update of the inter-limb connection is:
@@ V-A Problem Setup and Simulation Parameters
MuJoCo (https://github.com/deepmind/mujoco) version 2.1.0 is adopted in this paper for the quadruped robot simulation. mujoco-py [32], version 2.1.2.14, is used to call the MuJoCo functioning through Python APIs. The MJCF model of the A1 robot is obtained from the manufacturer’s repository (https://github.com/unitreerobotics/unitree_mujoco). The simulations are run on a bare metal computer with one Intel(R) Core(TM) i9-13900K CPU, 64GBytes RAM and the Ubuntu 22.04.2 LTS operating system.
@@
The spiking neuron models are implemented in Python in a discrete manner, where the simulation frequency is 1000Hz. The structure of the algorithm in a single time step can be split into three distinct steps:
@@ V-B Training Session and Training Progress
As the reward-modulated STDP reinforces the synapse weight when concurrent firing provides a positive reward, the synapse weights will continuously change even when a good gait allows the robot to run smoothly. Therefore the learning rate should be dependent on the learning progress, in order to halt the learning when a well-performing gait has been captured. The training progress is defined as:
@@
At the beginning of the entire training, the robot has a large possibility to fall to the ground. The training goal at the beginning thus is learning not to fall. Progressing further, at the stage where the robot can run continuously for 7 to 8 seconds, the learning goal becomes maximizing the rewards. The beginning 2 seconds of the session, at this time, negatively contributes to the learning goal because of the zero resetting velocity. Consequently, we designed a dynamic learning period instead of entire-session learning. The starting time step of the learning period in any session is:
@@
Figure 3: Ablation studies. a, The trained inter-limb connection weights. b, Typical thigh extensor activities of the four limbs after training. Upper panel: limb collaboration is absent when astrocyte regulation is removed. Lower panel: diagonal-synchronization emerges under astrocyte regulation. c, The average reward and the displacement in $x$ direction at the final time step of each session. The light color plots represent the raw value, and the dark lines represent the sliding window average of the raw value, with window size 5.
@@ V-C Gait Emergence and Ablation Study
The entire training process can be divided into multiple training sessions. For details about training sessions, please check the Methods section. At the beginning of the training, all the synapses of the inter-limb connection have zero values. The limbs hence swing back and forth individually with random and time-variant phase differences but similar frequencies. In the process of such random phase difference generation, multiple types of collaboration among the limbs will temporally emerge. For example, trotting, pacing or bounding can be seen for short time periods, typically much shorter than the maximum session length, at the beginning stages of learning. Based on our observations in the performed experiments, such random swings usually lead to a fall or non-continuous slow running. Among all the emerged gaits, the optimal gait produces maximum $x$-directional speed and stable attitude of the torso. Nevertheless, the non-optimal gaits, generating slow locomotion, will also result in a positive reward, because of the large reward coefficient of the $x$-directional speed. The effect of the optimal and the non-optimal gaits all positively contribute to the inter-limb connection, through the STDP learning algorithm. Such always-positive learning causes saturated inter-limb synapse strengths at the pre-defined strength upper bound. Hence, there is a strong need to build inhibitory inter-limb connection weights which is performed by the astrocytes through the release of adenosine to the motor neurons.
@@
Figure 3 shows the homeostatic contribution of the astrocytes. Removing the astrocyte regulation implies that the second term on the right hand side of Equation 13 is removed. In Figure 3a, the left panel shows the trained inter-limb connection weights without the astrocyte model. All the synapses are trained into excitatory connections with saturated weight $w_{\mathrm{max}}$. In the right panel, it can be observed that the connection from one limb to another limb always contains two excitatory and two inhibitory synapses. For example, the connections from FR limb to FL limb have two inhibitory synapses in the diagonal positions and two excitatory connections in the anti-diagonal positions, which implies that the extensor of FR inhibits the extensor of FL (similar discussions are valid for the flexors). The extensor of FR excites the flexor of FL and vice versa. Such a combination produces a complementary firing pattern of the FR and the FL. In fact, the right panel of Figure 3a indicates a diagonal-synchronizing gait: FR-RL synchronized and FL-RR synchronized. The two groups are complementary to each other. This diagonal-synchronizing gait is usually called trotting. In summary, the astrocyte can filter out the effect of non-optimal gaits. Figure 3b more obviously illustrates the trotting gait. In the upper panel, the all-excitatory inter-limb connection brings unstable locomotion where the phase difference among the limbs is unstructured. Under current thigh and calf limit positions, $I_{\mathrm{background}}$, $I_{\mathrm{thigh}}$ and $I_{\mathrm{calf}}$, only trotting gait will emerge. For a real quadruped animal, for example, a cat, the muscle tension and CPG rhythmic pattern frequency increase along with the descending modulating signal from the brainstem. Therefore, for the emergence of other gaits, the above-mentioned parameters need to be specifically tuned for our robotic application scenario. In Figure 3c, the average rewards in each session, the final $x$ coordinate and the session lengths are plotted. The final $x$ coordinate evaluates the distance the robot has moved as the reset heading direction is positive $x$ direction. The reward and moving distance in the case without astrocyte modulation fluctuate around a constant and do not significantly improve after training. In contrast, including astrocyte regulation improves the average reward and locomotion speed, because of the emergence of the optimal gait. More specifically, the average x-directional speed of the last 20 sessions is $1.17\,\mathrm{m/s}$, which is close to the trotting gate speed manually set in prior literature [9]. A session length shorter than 10 seconds implies a fall of the robot in the session. Involving astrocyte regulation significantly decreased the probability of falling during running.

### v3

@@ III-B2 Inter-limb Connection
Recent works have elaborated that there are excitatory and inhibitory synapses from the motor neurons for one limb to the motor neurons for another limb [35]. In this paper, the phase difference among the limbs is modulated by the inter-limb excitatory/inhibitory connections. As shown in Figure 1d, the thigh muscles from one limb send axons to all the muscles in all the other limbs. For example, the motor neurons of the thigh extensor in the front-left limb send connections to the thigh flexors and the extensors in front-right, rear-left and rear-right limbs. The calf motor neurons are not involved in the inter-limb connection. The source motor neuron pool and the target motor neuron pool are fully connected, but the entire connection shares a single synapse strength value. Such a connection could be expressed as a table of size 8 by 8, shown in Figure 2a, where the value in row $i$ and column $j$ represents the synapse weight of the connection projected from muscle $i$ to muscle $j$. The inter-limb connection is trained by reward-modulated Spike Timing Dependent Plasticity (STDP), including the necessary modulation of astrocytes, which will be introduced in a later section. In the next subsection, the cytoplasmic signaling pathway will be introduced. The computational detail regarding the inhibitory effect of astrocyte output on inter-limb connection will be illustrated in Section IV.
III-C Astrocyte Regulation
The astrocyte is reported to play a homeostatic role in spinal cord locomotion control circuitry [3, 5, 6]. Specifically, the astrocytes, when activated, release signaling molecules that suppress over-excited neural groups to prevent energy waste. The spiking activity of neighboring neurons is a trigger of the cytoplasmic calcium activity of astrocytes. In general, first, specific receptors on the plasma membrane of the astrocyte sense the product of the activated neighboring neurons, such as 2-arachidonyl glycerol (2-AG) or glutamate. Next, a sequence of interactions between $\mathrm{Ca}^{2+}$, inositol trisphosphate ($\mathrm{IP}_{3}$) and $\mathrm{IP}_{3}$ receptor responds to the activated membrane receptors. Finally, multiple types of signaling molecules are released from the astrocytes, such as glutamate or ATP, following the $\mathrm{Ca}^{2+}$ or $\mathrm{IP}_{3}$ activity.
Astrocytes are diverse across the entire nervous system. For the locomotion control system, a previous work suggested that Endocannabinoid receptor $\mathrm{CB}_{1}$ is expressed by the astrocytes in multiple brain regions and the dorsal horn of the spinal cord [43]. In this paper, we regard 2-AG, a type of endocannabinoid, as the effective neurotransmitter signaling from motor neurons to the astrocytes. The motor neurons release a constant amount of 2-AG every time there is an action potential. The concentration of 2-AG decays exponentially when no motor neuron fires:
For the astrocyte-to-neuron signaling pathway, it is well established that ATP is the direct product of astrocytes that suppresses the locomotion activity [3, 6, 4]. Nevertheless, the mathematical model for the relationship between the output signals of the astrocyte and the $\mathrm{Ca}^{2+}$ and $\mathrm{IP}_{3}$ activity in locomotion circuitry has not been detailed in prior works. Also, the detailed intra-astrocyte dynamics is absent in previous works on astrocyte-modulated locomotion. Instead, these works concentrate on the communication between astrocytes and the surrounding neurons [3, 6]. Therefore, a computational model for intra-astrocyte activity and ATP production needs to be adopted. A prior study [44] proposed a single-astrocyte model whose output is extracellular ATP. However, according to the paper, the evidence supporting the relationship between $\mathrm{IP}_{3}$ and ATP release was not concrete [44]. Chen et al. also state that the calcium activity is believed to be the direct cause of the rise in extracellular ATP concentration [6]. Further, another work proposed a data-driven mathematical model of astrocyte, which, nevertheless, concentrates on the propagation of calcium activity in an astrocyte network [45]. As a result, adopting a customized model for describing the intra-astrocyte activity and the astrocyte-ATP relationship is inevitable. Here, in the absence of backing of in vivo experiments on astrocyte modulation in the spinal cord, we use the Li-Rinzel model for intra-astrocyte dynamics and use the unit ATP release model for the ATP production from astrocytes, where the ATP release directly depends on the astrocyte cytoplasmic $\mathrm{Ca}^{2+}$ activity. This is consistent with results reported in Refs. [3, 6] and at the same time minimizes the number of redundant assumptions.
The Li-Rinzel model [46], which describes the general cytoplasmic $\mathrm{Ca}^{2+}$-$\mathrm{IP}_{3}$ dynamics, is the basis of many following astrocyte dynamics models [47]. In the Li-Rinzel model, the activation of $\mathrm{CB}_{1}$ receptor results in $\mathrm{IP}_{3}$ release in the astrocyte cytoplasm. $\mathrm{IP}_{3}$ further regulates the release of $\mathrm{Ca}^{2+}$ from the Endoplasmic Reticulum (ER) [24]. The transportation of $\mathrm{Ca}^{2+}$ between the cytoplasm and ER can be divided into three components: the channel current from ER to the cytoplasm, $J_{\mathrm{chan}}$; the passive leakage current from the ER to the cytoplasm, $J_{\mathrm{leak}}$; the active pumping current from the cytoplasm to ER, $J_{\mathrm{pump}}$. In summary:
The $\mathrm{Ca}^{2+}]$ here is the concentration of cytoplasmic $\mathrm{Ca}^{2+}$. Readers seeking a detailed computational model are directed to [24].
Chen et al. indicate that the extracellular ATP rises following the calcium activity in astrocytes [6]. The extracellular ATP, once released, is metabolized into adenosine immediately [3, 4, 6]. The adenosine plays an inhibitory role in the modulation of locomotion in the spinal cord [3, 6]. Multiple works have reported that the adenosine receptor on the membrane of the locomotion-related interneurons is a part of the inhibitory signaling pathway, but the specific type of receptor varies among different reports. Ref. [4] suggests that adenosine activates the $\text{A}_{1}$ receptor of the locomotion-related neurons in mice, while another report demonstrates that $\text{A}_{2}$ is the target of adenosine in swimming behavior modulation in zebrafish [6]. As ATP is always metabolized into adenosine rapidly [3, 4, 6], and the time for degeneration from ATP to adenosine can be neglected, we simplify the astrocyte-ATP-adenosine dynamics into astrocyte-adenosine dynamics for reducing the complexity of the computational model. The product of ATP metabolism, adenosine, is regarded as the direct output of the astrocyte. It is also assumed that one unit of adenosine $r_{\mathrm{ADO}}$ is released every time when the calcium concentration exceeds a threshold. Then, adenosine diffuses in the extracellular space and experiences exponential decay.
Here, $\mathrm{ADO}]$ is the concentration of adenosine in the nearby region of the astrocyte, $\tau_{\mathrm{ADO}}$ is the adenosine decay time constant, $r_{\mathrm{ADO}}$ is the production rate of adenosine, $t_{\mathrm{Ca}^{2+},\mathrm{ADO\_prev}}$ is the time instant when $\mathrm{Ca}^{2+}]$ is above the threshold satisfying the condition that the time from the last adenosine release is not shorter than the adenosine refractory period. The adenosine received by the motor neurons negatively contributes to the synapse strength sent to this motor neuron pool. Figure 2b illustrates the negative feedback relationship between the motor neuron pool and the adenosine.
@@ IV Gait Learning Algorithm
Previous works have discussed the reward-modulated STDP for encouraging actions through a delayed reward signal [48]. The basic idea of gait search is to reward the gait, i.e., the phase difference of the bursts among all the motor neuron pools, providing a larger forward speed and more stable torso attitude. The effect of different collaborations of the muscles is usually delayed, i.e., a good gait leads to a large forward speed, but the speeding up procedure still costs a duration of time. Hence, reward-modulated STDP is necessary for gait search. The reward function is given by:
@@
where, $\mathrm{STDP}_{x,y}(t)$ is the STDP signal from motor neuron pool $x$ to motor neuron pool $y$. $\tau_{\mathrm{STDP}}$ is the STDP signal decay time constant. $t_{y\mathrm{\_spike}}$ and $t_{x\mathrm{\_spike}}$ are the times of firing of motor neuron pool $y$ and $x$. $u_{x}(t)$ and $u_{y}(t)$ are the spike traces of motor neuron pool $x$ and $y$ at time $t$. $\tau_{\mathrm{trace}}$ is the spike trace decay time constant. $\eta_{\mathrm{negative\_relative}}$ is the relative negative learning rate, which controls the contribution of negative STDP signal in learning. $\eta_{\mathrm{eff}}(t)$ is the effective learning rate, which is the product of training progress $\mathrm{Progress}(t)$ and learning rate $\eta$. The definition of training progress can be found in the Experiments section. $r_{\mathrm{eff}}(t)$ is the effective reward for STDP learning. $\overline{r_{(t-0.1s,t)}}$ is the sliding window average of the reward function from 0.1 seconds before to the current time. $c_{\mathrm{average}}$ is the averaged reward coefficient. By subtracting a fraction of $\overline{r_{(t-0.1s,t)}}$, the system encourages the concurrent firing of the muscles which provides a higher reward than the average. It also discourages the concurrent firing of the muscles leading to the falling of the robot, at which time the angular velocities of the torso are large. $\zeta_{w_{x,y}}(t)$ is a weight constraint factor preventing the weight from being out of the range of $w_{\mathrm{min}},w_{\mathrm{max}}]$.
@@
It is worth noting here that the STDP learning algorithm is usually applied in situations where excitatory synapses play a significant role [49]. The learning rate for post-synaptic firing is usually much larger than its counterpart for pre-synaptic firing. As a result, the learning of inhibitory synapses, which are critical in many neural systems like ours, is largely unexplored. According to the previous section, adenosine released by the astrocytes is sensed by the neighboring neurons. For simplifying the computational model, the detailed interaction between the adenosine and motor neurons is neglected. It is assumed that the adenosine generates an inhibitory effect on the input synapses of the motor neurons. i.e., the adenosine reduces the weight of the inter-limb synapses from other motor neurons to the corresponding motor neuron pool, as illustrated in Figure 2b. Therefore, at every time step, the astrocyte-induced weight update from motor neuron pool $x$ to motor neuron pool $y$ is:
@@
where, $\eta_{\mathrm{ADO}}$ is the efficacy of adenosine in building inhibitory synapses, $\mathrm{Progress}(t)$ is the training progress, $\mathrm{ADO}_{y}]$ is the adenosine concentration at motor neuron pool $y$, $\zeta_{w_{x,y}}(t)$ is the weight constraint factor. The complete weight update of the inter-limb connection is:
@@
For discussion on the computational modeling of the proposed system, please refer to the section Computational Implementation of the Dynamical System in the supplementary material.
@@ V-A Problem Setup and Simulation Parameters
The experiments are performed using a computational simulation framework containing the quadruped robot physics simulation and the discrete simulation of the SNN-based CPG. MuJoCo (https://github.com/deepmind/mujoco) version 2.1.0 is adopted in this paper for the quadruped robot simulation. mujoco-py [50], version 2.1.2.14, is used to call the MuJoCo functions through Python APIs. The MJCF model of the A1 robot is obtained from the manufacturer’s repository (https://github.com/unitreerobotics/unitree_mujoco). The simulations are run on a bare metal computer with one Intel(R) Core(TM) i9-13900K CPU, 64GBytes RAM and the Ubuntu 22.04.2 LTS operating system. To simulate the dynamics of the SNN, i.e., the neuron/astrocyte models and the synaptic learning rules, we implemented the model dynamics in Python in a discrete manner, without utilizing any external SNN library. The simulation frequency is 1000 Hz. The entire codebase of the SNN implementation can be found at https://github.com/NeuroCompLab-psu/Astrocyte_Regulated_Gait_Search.git.
@@
The structure of the algorithm in a single time step can be split into three distinct steps:
@@ V-B Training Session and Training Progress
As the reward-modulated STDP reinforces the synapse weight when concurrent firing provides a positive reward, the synapse weights will continuously change even when a good gait allows the robot to run smoothly. Therefore, the learning rate should be dependent on the learning progress, in order to halt the learning when a well-performing gait has been captured. The training progress is defined as:
@@
At the beginning of the entire training, the robot has a high possibility to fall to the ground. The training goal at the beginning, therefore, is learning not to fall. Progressing further, at the stage where the robot can run continuously for 7 to 8 seconds, the learning goal becomes maximizing the rewards. The beginning 2 seconds of the session, at this time, negatively contribute to the learning goal because of the zero resetting velocity. Consequently, we designed a dynamic learning period instead of entire-session learning. The starting time step of the learning period in any session is:
@@
Figure 3: Ablation studies. a, Typical trained inter-limb connection weights. b, Typical thigh extensor activities of the four limbs after training. Upper panel: limb collaboration is absent when astrocyte regulation is removed. Lower panel: diagonal-synchronization emerges under astrocyte regulation. c, The statistics of the average reward, the displacement in $x$ direction at the final time step of each session, and session length in the training process, where the number of simulation trails is 20. The dark line represents the mean value, and the shaded region represents the interval of mean $\pm$ standard deviation.
@@ V-C Gait Emergence and Ablation Study
The entire training process can be divided into multiple training sessions. For details about training sessions, please check the Methods section. At the beginning of the training, all the synapses of the inter-limb connection have zero values. The limbs hence swing back and forth individually with random and time-variant phase differences but similar frequencies. In the process of such random phase difference generation, multiple types of collaboration among the limbs will temporally emerge. For example, trotting, pacing or bounding can be seen for short time periods, typically much shorter than the maximum session length, at the beginning stages of learning. Based on our observations in the performed experiments, such random swings usually lead to a fall or non-continuous slow running. Among all the emerged gaits, the optimal gait produces maximum $x$-directional speed and stable attitude of the torso. Nevertheless, the non-optimal gaits, generating slow locomotion, will also result in a positive reward, because of the large reward coefficient of the $x$-directional speed. The effect of the optimal and the non-optimal gaits all positively contribute to the inter-limb connections, through the STDP learning algorithm. Such always-positive learning causes saturated inter-limb synapse strengths at the pre-defined strength upper bound. Hence, there is a strong need to build inhibitory inter-limb connection weights, which is performed by the astrocytes through the release of adenosine to the motor neurons.
@@
Figure 3 shows the homeostatic contribution of the astrocytes. Removing the astrocyte regulation implies that the second term on the right-hand side of Equation 13 is removed. In Figure 3a, the left panel shows the trained inter-limb connection weights without the astrocyte model. All the synapses are trained into excitatory connections with saturated weight $w_{\mathrm{max}}$. In the right panel, it can be observed that the connection from one limb to another limb always contains two excitatory and two inhibitory synapses. For example, the connections from FR limb to FL limb have two inhibitory synapses in the diagonal positions and two excitatory connections in the anti-diagonal positions, which implies that the extensor of FR inhibits the extensor of FL (similar discussions are valid for the flexors). The extensor of FR excites the flexor of FL and vice versa. Such a combination produces a complementary firing pattern of the FR and the FL. In fact, the right panel of Figure 3a indicates a diagonal-synchronizing gait: FR-RL synchronized and FL-RR synchronized. The two groups are complementary to each other. This diagonal-synchronizing gait is usually called trotting. In summary, the astrocyte can filter out the effect of non-optimal gaits. Figure 3b more obviously illustrates the trotting gait. In the upper panel, the all-excitatory inter-limb connection brings unstable locomotion where the phase difference among the limbs is unstructured. Under current thigh and calf limit positions, $I_{\mathrm{background}}$, $I_{\mathrm{thigh}}$ and $I_{\mathrm{calf}}$, only trotting gait will emerge. For a real quadruped animal, for example, a cat, the muscle tension and CPG rhythmic pattern frequency increase along with the descending modulating signal from the brainstem. Therefore, for the emergence of other gaits, the above-mentioned parameters need to be specifically tuned for our robotic application scenario. Furthermore, this trotting gait is currently only applicable to flat ground. Adaptability to more complex terrain will be the scope of future work. In Figure 3c, the average rewards in each session, the final $x$ coordinate and the session lengths are plotted. The final $x$ coordinate evaluates the distance the robot has moved as the reset heading direction is positive $x$ direction. The reward and moving distance in the case without astrocyte modulation fluctuate around a constant and do not significantly improve after training. In contrast, including astrocyte regulation improves the average reward and locomotion speed, because of the emergence of the optimal gait. More specifically, the average x-directional speed of the last 20 sessions is $1.17\,\mathrm{m/s}$, which is close to the trotting gate speed manually set in prior literature [20]. Such a speed emerges naturally, and it is the result of specific CPG burst duration and motor torque hyper-parameters. A session length shorter than 10 seconds implies a fall of the robot in the session. Involving astrocyte regulation significantly decreased the probability of falling during running.